# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have led to significant successes across various applications, where the most noticeable is to a series of emerging capabilities, particularly in the areas of In-Context Learning (ICL) and Chain-of-Thought (CoT). To better understand and control model performance, many studies have begun investigating the underlying causes of these phenomena and their impact on task outcomes. However, existing explanatory frameworks predominantly focus on isolating and explaining ICL and CoT independently, leading to an incomplete understanding of their combined influence on model performance. To address this gap, we propose the Electronic Circuit Model (ECM), which provides a foundation for developing scalable, learnable policies and improving the management of AI-generated content. Specifically, ECM conceptualizes model behavior as an electronic circuit: ICL is represented as semantic magnetic field to providing an additional voltage following Faraday’s Law, while CoT is modeled as series resistors to constrain the model output performance following Ohm’s Law. Experimental results demonstrate that the ECM effectively predicts and explains LLM performance across a variety of prompting strategies. Furthermore, we apply ECM to advanced reasoning strategy optimization on a series of tasks, such as the International Olympiad in Informatics (IOI) and the International Mathematical Olympiad (IMO), achieving competitive performance that surpasses nearly 80% of top human competitors.
Key Words: Electronic Circuit Model, Large Language Model, In-Context Learning, Chain-of-Thought
@@ 1 Introduction
Large Language Models (LLMs) have made significant strides across various tasks, including code generation, knowledge-based question answering, and complex reasoning [46, 3, 29, 32]. Unlike earlier models that showed incremental improvements, LLMs exhibit emergent capabilities that fundamentally reshape problem-solving. At the core of these capabilities are two critical mechanisms: In-Context Learning (ICL) [2, 11, 31] and chain-of-Thought (CoT) reasoning [26, 41, 5], enabling LLMs to adapt dynamically to new challenges and achieve human-comparable learning and reasoning [2, 27, 1]. Specifically, ICL triggers LLMs to generalize information based on task-specific context provided in the input, without requiring parameter updates. In parallel, CoT reasoning allows LLMs to approach complex problems stepwise, enhancing both performance and interpretability. Together, these mechanisms empower LLMs to serve as versatile problem solvers in critical fields such as healthcare and education [14, 25, 16, 18].
Despite their transformative potential, the mechanism of LLMs’ emergent capabilities remains poorly understood. Current research on ICL has uncovered preliminary evidence of how models encode and process contextual information, with studies suggesting latent representations that approximate forms of task-specific meta-learning [38, 36] or topic models [39]. Similarly, CoT research has begun to map the reasoning pathways LLMs employ, providing insights into their demonstration alignment with human cognitive heuristics [24, 37]. Others seek to interpret emergent behaviors using macroscopic mathematical models based on experimental analysis [13, 5]. In practice, ICL and CoT are often integrated, allowing models to adapt to context and solve problems step-by-step [17, 15, 30, 47, 6]. To put it figuratively, combining ICL and CoT is like planting a tree: ICL provides the rich “soil” in which the model can absorb and integrate contextual knowledge, while CoT guides its “branches” to grow step by step in a clear, structured direction. Together, they create a natural, human-like reasoning process that’s both flexible and transparent, enabling models to tackle complex tasks with greater depth and precision. Unfortunately, in the current literature, a unified understanding of the combined effect of ICL and CoT remains limited, which seriously affects correct performance interpretation and controllable result prediction and further constrains the development of principled methodologies for LLM design and deployment.
Figure 1: Glossary of Terms and Schematic of Electronic Circuit Model (ECM).
In-Context Learning (ICL) is conceptualized as a (Semantic) Magnetic Field. Demonstrations in ICL effectively provide an additional voltage by applying a magnetic field change aligned with the input query direction. Chain-of-Thought (CoT) process is analogized as series resistors in circuit, where the reasoning difficulty in CoT are divided into a sequence of sub-difficulties, analogous to resistors in a circuit that collectively increase overall resistor. Building on this, we introduce the concept of model inherent capabilities, represented as the power supply of the circuit to provide basic voltage. Model performance can then be quantified as the circuit’s output power, enabling more precise predictions of empirical accuracy. In addition, in this article, “strength” denotes the value of semantic magnetic fields, “resistance” denotes the value of resistor, and “voltage” denotes the value of power supply.
To bridge this gap, we introduce the Electronic Circuit model (ECM), a unified theoretical paradigm that reimagines model performance through the lens of physical circuit systems. Specifically, as shown in Figure 1, we try to conceptualize ICL capabilities as semantic magnetic fields, driving additional voltage gains that correspond to enhanced information absorption, while CoT challenges are modeled as series resistors, introducing predictable constraints akin to circuit losses. Based on these hypotheses, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning process. By rigorously aligning theoretical circuit power predictions with empirical accuracy results across 3 prompting strategies, 4 task categories, and 13 different LLMs. The ECM successfully establishes a unified quantitative paradigm for predicting and optimizing the behavior of LLMs.
Moreover, based on the introduction of ECM, we further explain over 10 existing prompt optimization strategies from the perspective of optimizing the magnetic field and optimizing the resistors. And given these perspectives, some phenomena that have not been successfully explained by previous works and more excellent prompt strategies are discovered based on ECM. Notably, our results show that combined with these approaches has achieved breakthroughs in complex domains: LLMs guided by ECM-optimized strategies have outperformed over 80% of human competitors in the International Olympiad in Informatics and the International Mathematical Olympiad. Furthermore, in subjective and exploratory contexts such as academic research development, ECM-optimized strategies have driven significant gains, achieving a minimum 10% improvement in overall human scores.
@@
  • To the best of our knowledge, we are taking the first meaningful step to introduce the Electronic Circuit Model (ECM), effectively offering a unified theoretical paradigm to reformulate model performance through the lens of circuit systems.
  • In this framework, we discover Faraday’s Law of ICL and Ohm’s Law of CoT in LLMs’ reasoning, and further utilize ECM to effectively predict the empirical accuracy and explain the effectiveness of 10 existing strategies from the view of optimizing the magnetic field and optimizing the resistor.
  • We utilize ECM to explain some open phenomena and to propose more excellent prompt strategies based on ECM. Combined with these approaches, we have achieved breakthroughs in mathematical and code competition and academic research development scenarios. Further, we empirically demonstrate that applying ECM-driven optimizations to LLMs leads to substantial performance gains.
2 Theoretical Model
This theoretical model shows that model performance ($P_{\text{out}}$) relies not only on basic voltage from model inherent capabilities ($\mathcal{E}_{\text{model}}$) but also on external factors like reasoning difficulty ($R_{CoT}$) and additional voltage from extra demonstrations in ICL ($\mathcal{E}_{ICL}$), which are described as follows in detail:
@@ 2.1 Model Inherent Capabilities as Power Supply
In our model, we draw an analogy by conceptualizing LLM’s inherent capabilities as a power supply that drives computational processes and reasoning tasks (Figure 1). The voltage value $\mathcal{E}_{\text{model}}$ represents the strength of the inherent capability, governing the LLM’s capacity to execute a range of tasks. Just as the voltage in an electrical circuit determines the power available to drive various components, the internal capabilities govern its ability to manage increasingly complex inputs.
2.2 Faraday’s Law of In-Context Learning
Inspired by Wang et al. [38], the process of ICL can be conceptualized as the interaction of LLM with contextual data, where it absorbs, retains, and eventually releases semantic information flow. This dynamic flow of information resembles the changing behavior of a magnetic field as its initial strength $\Phi^{B}_{0}$ decreases to zero. Building on this analogy, we propose Faraday’s Law for the ICL paradigm, where the rate of change in the “magnetic field" further induces an extra voltage:
$\mathcal{E}_{\text{ICL}}=-\frac{d\Phi^{B}(t)}{dt}=\lambda\Phi_{0}^{B},$ (1)
where $\lambda$ represents the uniform rate of decrease in magnetic field intensity. This perspective enables a deeper quantitative understanding of ICL through the concept of a “semantic magnetic field," as illustrated in Figure 1.
2.3 Ohm’s Law of Chain-of-Thought
In a parallel vein, Chen et al. [5] have discovered a combination law for LLM reasoning, which aligns with the series formula of electrical resistors after transformation (See in Appendix A). Inspired by this, we conceptualize the difficulty of CoT reasoning as analogous to series circuits. In this framework, each reasoning process introduces a distinct resistor ($R_{\text{CoT}}=\sum_{i}R_{i}$), compounding the task’s overall difficulty ($R_{i}$). As illustrated in the right panel of Figure 1, Ohm’s law for CoT can be expressed as:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (2)
where $R_{0}$ represents the resistance of an output resistor with static resistance value in the circuit, which reflects the difficulty of reasoning conclusion. This formulation quantifies the cumulative effect of reasoning resistors, providing a modular quantitative perspective on cognitive task complexity.
@@ 2.4 Model Performance as Output Power of Bulb
Extending the conceptual model, we equate model performance to the power output of a light bulb in an electronic circuit. Task execution efficiency depends on a nonlinear interaction of internal and external factors, rather than solely on the model’s intrinsic capabilities. The output power, $P_{\text{out}}$, is expressed as:
$P_{\text{out}}=I^{2}_{model}R_{0}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{\text{CoT}}+R_{0})^{2}}.$ (3)

### v2

# Electronic Circuit Principles of Large Language Models

@@ Abstract
Large language models (LLMs) such as DeepSeek-R1 have achieved remarkable performance across diverse reasoning tasks. To uncover the principles that govern their behaviour, we introduce the Electronic Circuit Principles (ECP), which maps inference-time learning (ITL) onto a semantic electromotive force and inference-time reasoning (ITR) onto a resistive network governed by Ohm’s and Faraday’s laws. This circuit-based modelling yields closed-form predictions of task performance and reveals how modular prompt components interact to shape accuracy. We validated ECP on 70,000 samples spanning 350 reasoning tasks and 9 advanced LLMs, observing a about 60% improvement in Pearson correlation relative to the conventional inference-time scaling law. Moreover, ECP explains the efficacy of 15 established prompting strategies and directs the development of new modular interventions that exceed the median score of the top 80% of participants in both the International Olympiad in Informatics and the International Mathematical Olympiad. By grounding LLM reasoning in electronic-circuit principles, ECP provides a rigorous framework for predicting performance and optimising modular components.
Key Words: Electronic Circuit Principles, Large Language Model, Inference-time Learning, Inference-time Reasoning
Figure 1: The Introduction of Inference-Time Learning (ITL) and Inference-Time Reasoning (ITR), and Comparison of Micro- and Macro-Circuits.
a, Introduction to ITL, which pairs queries with solution in input context as references, and to ITR, which provides step-by-step reasoning in model-generated output. b, Electronic Circuit Principles on Macro-circuit analyses: Traditional micro-circuits capture basical linguistic functions but miss higher-order reasoning principles. The micro-circuit structure follows Rai et al. [35].
@@ 1 Introduction
Large Language Models (LLMs), like DeepSeek-R1 [14] and OpenAI-o1 [17], have recently achieved unprecedented performance in code generation, knowledge-based question answering and complex reasoning tasks [6, 25, 34]. Unlike conventional software, whose behavior is explicitly encoded by human, LLMs exhibit emergent, stochastic dynamics that defy straightforward performance modeling. As a result, users often cannot anticipate when a model will succeed or fail before inference and generation, creating a “Schrödinger’s cat” scenario in which reliability only becomes apparent post hoc.
At inference, as shown in Figure 1a, two complementary behaviors drive LLM adaptability [6]: inference-time learning (ITL) [2, 11, 33] and Inference-Time Reasoning (ITR) [30, 43, 4]. ITL, often termed in-context learning, enables parameter‐free generalization from examples supplied in the prompt. By contrast, ITR, frequently implemented via chain-of-thought prompting, decomposes complex tasks into interpretable reasoning steps, enhancing both accuracy and transparency. These paradigms have underpinned rapid advances in domains ranging from clinical decision support to personalized education [15, 29, 20, 22], yet they also obscure the behind rationales governing performance, impeding systematic understanding and theory-guided refinement of reasoning strategies from higher-order perspectives.
To address this gap, we present the Electronic Circuit Principle (ECP), which maps ITL and ITR onto analogues of classical circuit elements. As illustrated in Fig. 1b, in this principle, ITL acts as an inference-time‐varying electromotive force, enabling demonstration incorporation much like induced voltage. Meanwhile, ITR maps onto an electrical network in which each reasoning subtask contributes series resistance, such that the total resistance quantifies cumulative cognitive load. By formalizing Faraday’s and Ohm’s laws of circuits, ECP furnishes a principled framework that both predicts LLM performance and guides modular optimization of reasoning components.
We validate this approach on nine widely used LLMs across 350 tasks (over 70,000 samples), achieving a Spearman correlation > 0.7 between ECP‐based predictions and empirical accuracy. This represents about 60% correlation improvement over current inference-time scaling law methods that account for post-inference results, highlighting the enhanced predictive capacity of ECP. Moreover, ECP‐inspired adjustments, treating demonstration strength optimization as “voltage” improvement and reasoning strategy optimization as “resistance” reduction, consistently improve model performance based on theoretical insights.
Notably, LLMs employing ECP-optimized strategies surpassed 80% of human competitors in the International Olympiad in Informatics (IOI) and the International Mathematical Olympiad (IMO) based on much weaker LLMs. Furthermore, in subjective and exploratory domains such as academic research, ECP-optimized strategies contributed to a minimum 10% improvement in human performance metrics. These findings demonstrate the utility of the Electronic Circuit principle for targeted modular optimization.
@@
  • Uncovering Macro-level Principles in LLMs: We introduce the Electronic Circuit Principle (ECP), a pioneering framework that redefines model performance through the lens of macro-circuit, providing a unified theoretical paradigm for quantitatively predicting and optimizing LLM performance.
  • Pioneering Accurate Performance Prediction: By drawing analogies to Faraday’s Law of Induction and Ohm’s Law, we offer new insights into LLM reasoning mechanisms, enabling precise performance predictions with a Spearman correlation exceeding 70% from 9 widely-used LLMs across more than 350 tasks.
  • Guiding Modular Performance Optimization: Leveraging ECP, we interpret existing optimization strategies and propose novel optimization strategies informed by ECP theory. These strategies have led to significant breakthroughs in various fields, from mathematical and programming competitions to cutting-edge academic research.
Figure 2: Glossary of Terms and Schematic of Electronic Circuit Principles (ECP).
Inference-time learning (ITL) is depicted as a dynamic semantic magnetic field whose strength rises when demonstrations accompany the prompt, adding potential that steers token selection. Inference-time reasoning (ITR) is modelled as a series circuit: each reasoning sub-task acts as a resistor, and their sum gives the total cognitive load. The LLMs’ intrinsic capacity supplies the driving voltage, and empirical accuracy corresponds to the power delivered across the load. “Strength”, “Resistance” and “Voltage” map to field intensity, sub-task difficulty and LLM capability, respectively, offering a compact framework to quantify and compare LLM behaviour.
2 Electronic Circuit Principles
Electronic Circuit Principles show that model performance ($P_{\text{out}}$) relies not only on basic voltage from model inherent capabilities ($\mathcal{E}_{\text{model}}$) but also on external factors like reasoning difficulty ($R_{ITR}$) and additional voltage from extra demonstrations in ITL ($\mathcal{E}_{ITL}$), which are described as follows in detail:
@@ 2.1 Model Inherent Capabilities as Power Supply
In ECP, we draw an analogy by conceptualizing LLM’s inherent capabilities as a power supply that drives computational processes and reasoning tasks (see in Fig. 2). The voltage value $\mathcal{E}_{\text{model}}$ represents the strength of the inherent capability, governing the LLM’s capacity to execute a range of tasks. Just as the voltage in an electrical circuit determines the power available to drive various components, the internal capabilities govern its ability to manage increasingly complex inputs.
2.2 Faraday’s Law of Inference-Time Learning
Inspired by Wang et al. [40], the process of ITL can be conceptualized as the interaction of LLM and contextual data, whereby the model absorbs, retains, and ultimately releases semantic information. This dynamic flow of information is analogous to the behavior of a magnetic field, where its initial strength, $\Phi^{B}_{0}$, diminishes to zero over time. Based on this analogy, as illustrated in Fig. 2 (left), we introduce Faraday’s Law within the context of ITL, where the rate of change in the "magnetic field" induces a corresponding voltage. Specifically, we define the voltage as:
$\mathcal{E}_{\text{ITL}}=-\frac{d\Phi^{B}(t)}{dt}=\lambda\Phi_{0}^{B},$ (1)
where $\lambda$ denotes the uniform rate of magnetic field intensity decay. This framework provides a quantitative approach to understanding ITL by the concept of a “semantic magnetic field”.
2.3 Ohm’s Law of Inference-Time Reasoning
Similarly, Chen et al. [4] have derived a combination law for reasoning within LLM, which parallels the series formula for electrical resistors (See Methods). Drawing on this, we propose a principle for the difficulty of ITR that is analogous to series circuits. In this principle, each reasoning process introduces a distinct resistor ($R_{\text{ITR}}=\sum_{i}R_{i}$), compounding the task’s difficulty ($R_{i}$). As illustrated in Fig. 2 (right), Ohm’s law for ITR follows:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}}{R_{\text{ITR}}+R_{0}},$ (2)
where $R_{0}$ represents the resistance of an output resistor with static resistance value in the circuit, which reflects the difficulty of reasoning conclusion. This formulation quantifies the cumulative effect of reasoning resistors, providing a modular quantitative perspective on task complexity.
@@ 2.4 Model Performance as Output Power of Bulb
Extending the ECP, we equate model performance to the power output of a light bulb in an electronic circuit. Task execution efficiency depends on a nonlinear interaction of internal and external factors, rather than solely on the model’s intrinsic capabilities. The output power, $P_{\text{out}}$, can be formally expressed as:
$P_{\text{out}}=I^{2}_{model}R_{0}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{\text{ITR}}+R_{0})^{2}}.$ (3)